OpenStudy (anonymous):

helps meh again ? w varies jointly as x and y and inversely as the square of z. If w = 280 when x = 30, y = 12, and z = 3, find w when x = 20, y = 10 and z = 2.

OpenStudy (anonymous):

i believe the answer is between 0 and a billion :P

OpenStudy (anonymous):

omg you are so smart

OpenStudy (anonymous):

oh stop i blushh :P

OpenStudy (anonymous):

oh stop it yew haha

OpenStudy (callisto):

w varies jointly as x and y and inversely as the square of z w(z^2)/xy = constant = 280(3^2) / [(30)(12)] = 7 w = 7(xy) / (z^2) = 7x20x10/2^2 = 1400/4 = can you work out the answer?

OpenStudy (anonymous):

Hahaa :)

OpenStudy (anonymous):

350?

OpenStudy (callisto):

I think so, do you understand the workings?

OpenStudy (anonymous):

kind of. enough.